[figure: 9.7 cm]

THE USE OF AC SUSCEPTIBILITY MEASUREMENTS IN THE INVESTIGATION OF AGGREGATION IN MAGNETIC FLUIDS

P.C. Fannin^{a}, S.W. Charles^{b}

aDepartment of Electronic & Electrical Engineering, Trinity College, Dublin 2, Ireland. bDepartment of Chemistry, University College of North Wales, Bangor LL57 2UW, UK.

Measurements of the complex, frequency dependent susceptibility, χ(ω) = 3Dχ′′(ω) − iχ′′(ω), for a number of magnetic fluids over the frequency range 10Hz to 1MHz are presented. Plots of the susceptibility components against log (f Hz) are shown to have approximate Debye-type profiles with the presence of Brownian relaxation being indicated by the frequency, f_{max}, of the maximum if the χ′′(ω) profiles. Corresponding calculations of particle hydrodynamic radius indicates the presence of aggregation. Fitting techniques are used to obtain a measure of the particle size distribution in the samples.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : P.C. Fannin

Address for correspondence : Department of Electronic & Electrical Engineering Trinity College Dublin 2, Ireland.

Email address : pfannin@tcd.ie